OpenStudy (anonymous):

Find the constant of proportionality. R is inversely proportional to the square of I. If I = 20, then R = 100.

OpenStudy (dumbcow):

R = k/I^2 100 = k/20^2 100 = k/400 40,000 = k

OpenStudy (anonymous):

SO THE ANSWER IS 40,000?